Determine the value of an investment in the amount of $5000 that matures in 4 years and pays 4.5% per year compounded as indicated.
How to use the anual formula for compound interest here?

Answer by lwsshak3(11628)

Determine the value of an investment in the amount of $5000 that matures in 4 years and pays 4.5% per year compounded as indicated.
How to use the annual formula for compound interest here?
***
compound interest formula: A=P(1+i)^n, P=initial investment, i=interest rate per period, n=number of compounding periods, A=amt after n-periods.
..
For given problem:
P=5000
i=4.5%
n=4
..
A=5000(1+.045)^4
A=5000(1.045)^4≈$5963
..
note: I use a modified version of the compound interest formula found in most textbooks.
In this case, the compounding period is one year.